# LightRetriever: A LLM-based Hybrid Retrieval Architecture with 1000× Faster Query Inference

### v3

@@ 4 Ablations
Table 3: Symmetry Ablation settings (A1-A3).
@@
A3 1-layer Transformers Full 1-layer Transformers Full
@@
Table 4: Ablations results of LightRetriever
Ablations / nDCG@10 BeIR CMTEB-R
Llama-1b 48.7 58.7
w/ Both-side Lightweight 34.9-13.8 40.1-18.6
w/ Encq use a MLP 37.5-11.2 41.3-17.4
Llama-8b 54.4 63.0
w/ 1st-layer on query 50.1-4.3 54.4-8.6
@@ A.2.1 Text retrieval
Text retrieval builds on top of the vector space model Salton [1975], which encodes query and document representation vectors, and performs relevance search based on vector similarities.
@@
Dense retrieval Karpukhin et al. [2020] trains PLM-based encoders to condense queries and passages into low-dimensional dense vectors, then performs relevance search based on the Maximum Inner Product Search (MIPS) Mussmann and Ermon [2016] algorithm. In recent years, dense retrievers, such as BGE Chen et al. [2024] and E5 Wang et al. [2024a], have gained popularity for strong retrieval abilities across different tasks Muennighoff et al. [2023], Thakur et al. [2021], languages Zhang et al. [2023], and context granulaties Williams et al. [2018], Nguyen et al. [2016], owing to diverse training data Reimers [2019], retrieval-customized pre-training Gao and Callan [2021], Xiao et al. [2022], Ma et al. [2024], improved negative-mining Xiong et al. [2021], and so on. Recent SOTA retrievers Wang et al. [2024b], BehnamGhader et al. [2024], Lee et al. [2024] start to utilize LLMs Jiang et al. [2023], Touvron et al. [2023] as backbone encoders. With larger parameters and pre-training data, LLM-based retrievers enable significantly powerful retrieval capacities and domain adaptation abilities. However, LLMs also incur heavy online workloads when serving as query encoders. Improving the serving efficiency of LLM-based retrievers is still left for exploration.
@@
Traditional statistical sparse retrieval algorithms (e.g., TF-IDF Salton and Buckley [1988], and BM25 Robertson et al. [1994]) do not involve learnable language models. Instead, they directly analyze lexical frequencies (e.g., TF) and importance (e.g., IDF) in a large vocabulary space. These lexicals are sparsely distributed and require building inverted indexes Luhn [1957] for feasible similarity metric calculation, such as Apache Lucene Bialecki et al. [2012] and Tantivy Quickwit Inc. [2024]. In recent years, sparse retrieval has started to work with pre-trained LMs (PLMs). BGE Chen et al. [2024], SPLADE Formal et al. [2021], SparTerm Bai et al. [2020], and SPARTA Zhao et al. [2021] use customized XLM-RoBERTa Conneau et al. [2020] or BERT Devlin et al. [2019] as backbone encoders to learn term frequencies in an end-to-end manner and directly perform impact search based on the dot product of sparse vectors. Because of sparsity, these works also search with inverted indexes. Recent works Zhuang et al. [2024] also try to incorporate LLMs with sparse retrieval. However, the online efficiency issue still exists.
@@
Hybrid retrieval is a technique that combines retrieval scores Wang et al. [2021] or ranks Cormack et al. [2009] from multiple retrieval systems, such as dense Karpukhin et al. [2020] / sparse Robertson et al. [1994], Formal et al. [2021] retrieval, and single Karpukhin et al. [2020] / multiple Khattab and Zaharia [2020], Chen et al. [2024] vector retrieval. It enables better retrieval abilities by interpolating results from multiple sources. Linear interpolation Wang et al. [2021] of dense and sparse scores is used in our work.
@@ A.2.2 Improving inference efficiency
Previous works have tried to distill small query encoders from large encoders for better query inference efficiency. KALE Campos et al. [2023] prunes the BERT-based query encoder and distills it during post-training. However, when the query encoder is pruned to one Transformers layer, its recall@20 on NQ Kwiatkowski et al. [2019] decreases by 25-30%. Wang and Lyu [2023] proposes asymmetric dual-encoders by pruning BERT layers or using smaller BERT models. Its experiments show that a 2-layer BERT query encoder retains 92.5% of the full-sized performances via distillation and proper student initialization. USTAD Kim et al. [2024] also proposes similar asymmetry and distillation strategies, which choose the 12-layer BERT as teachers and the 6-layer DistilBERT or 2-layer BERT-mini as students. Its experiments show that the distilled BERT-mini query encoder achieves 95-97% of the teacher’s performance because of better distillation.
However, these studies have primarily focused on small language models (SLMs), such as BERT Devlin et al. [2019], and overlooked feasible approaches for improving online inference efficiency of LLM-based retrievers, where such optimization is even more critical. Furthermore, these existing methods typically retain Transformer layers in the query encoders, which limits their efficiency gains. In addition, most of these works rely on SLMs and are trained on relatively narrow datasets, such as MS-MARCO Nguyen et al. [2016], thus unable to compete with LLM-based retrievers.
@@
Previous works have also been dedicated to reducing index sizes for smaller disk space usage. Matryoshka representation learning (MRL) Kusupati et al. [2022] supports adjusting the dense vector dimensions by directly taking the top-k dimension of the original vector as shrink vectors. It aligns and trains the corresponding multi-dimensional vectors in the end-to-end contrastive learning. For example, if $k\in\{64,128,256\}$, query-passage dense vectors with $\{64,128,256\}$-dims could be obtained together after one encoder forward operation. For LM-based sparse retrieval, the similarity scores are directly affected by the impacts of corresponding lexicals Formal et al. [2021]. Thus, it naturally supports reducing the index sizes by retaining Top-k terms/lexicals. To explore the potential of controlling index sizes, our work also explores shrinking index sizes with dense MRL and sparse Top-k in the Section A.3, which also shows strong generalization capacities.
@@ A.3 Additional results about controls of vector dimension and sparsity
Retrieval systems usually involve indexing millions of documents into search databases, which consumes large amounts of disk space, especially for dense vectors. For example, indexing 8.8M MS-MARCO passages Nguyen et al. [2016] with Llama-8b model consumes 134.92 GB for Faiss dense retrieval and 11.24 GB for Lucene sparse retrieval. Thus, it’s essential to explore the possibility of controlling index sizes by shrinking vectors. 1) For dense retrieval, we incorporate Matryoshka Representation Learning (MRL) Kusupati et al. [2022] with LightRetriever to support flexible controls of dense vector dimensions. Given a dense vector $v^{den}$ with dimension $H$, MRL derives dimension-shrunk vectors by simply cutting out the Top-k dimensions $v^{den}[:k]$, then applies the same training loss as the original vectors. MRL trains a collection of Top-k dimensions together, enabling flexible choosing of a proper k. 2) For sparse retrieval, the sparse vector directly learns the term impact (i.e., lexical frequency) within a large sparse dimension. Thus, LightRetriever supports controlling sparsity by directly taking the Top-k terms, where the larger impact terms matter more.
@@ A.4 Additional ablation about the effect of auxiliary KL loss
Previous works for BERT-based asymmetric retrievers Wang and Lyu [2023], Kim et al. [2024] emphasize the importance of alignment between asymmetric students and full symmetric teachers. To explore the effect of such alignment, besides contrastive loss as the main loss function, we also explore to adopt the KL loss as an auxiliary to align the asymmetric similarity scores $S^{student}=v_{q}^{den}\cdot v_{d}^{den}$ to full-sized symmetric similarity scores $S^{teacher}=Enc_{q}(q)\cdot Enc_{d}(d)$,
@@ A.5 Training datasets
Following the settings of previous LLM-based retrievers Wang et al. [2024b], BehnamGhader et al. [2024], Ma et al. [2025], our work uses 23 training sets pre-built by tDRO Ma et al. [2025] or Sentence Transformers Training Data Reimers [2019], including 20 English and 3 Chinese datasets. Detailed dataset information is listed in Table 6 as follows.
@@
Amazon Review (2018) Ni et al. [2019] English Amazon $999$ 0.01 0.58%
cMedQA2 Zhang et al. [2018] Chinese Chinese Medical 99936 1 5.84%
DuReader Qiu et al. [2022] Chinese Chinese Web Collections 86366 1 5.05%
Eli5 Fan et al. [2019] English Reddit 325390 0.05 0.95%
Fever Thorne et al. [2018] English Wikipedia QA 109808 0.8 5.13%
FiQA Maia et al. [2018] English Financial 5498 1 0.32%
GooQA Pairs Khashabi et al. [2021] English Web Collections 3012347 0.025 4.40%
HotpotQA Yang et al. [2018b] English Wikipedia QA 85000 0.5 2.48%
MSMARCO Nguyen et al. [2016] English Web Collections 502854 1 29.39%
NFCorpus Boteva et al. [2016] English Medical 2585 1 0.15%
NPR Lucas et al. [2023] English News 594376 0.05 1.74%
NQ Kwiatkowski et al. [2019] English Wikipedia QA 58800 1 3.44%
PQA Pairs Lewis et al. [2021] English Wikipedia QA 99999 0.5 2.92%
Quora Duplicates Triples Iyer et al. [2012] English Forum Duplicates 97011 0.5 2.83%
S2ORC Title-Abstract Lo et al. [2020] English Semantic Scholar 99998 0.5 2.92%
SciFact Wadden et al. [2020] English S2ORC 806 1 0.05%
SPECTER Cohan et al. [2020] English Semantic Scholar 136642 0.25 2.00%
StackExchange Duplicates (Title-Body) da Silva et al. [2018] English Forum Duplicates 250516 0.25 3.66%
E5 Synthetic BeastyZ [2024] English GPT Synthetic Data 224791 1 13.14%
T2Ranking Xie et al. [2023] Chinese Chinese Web Collections 200362 0.5 5.86%
Trivia Joshi et al. [2017] English Wikipedia QA 60370 0.5 1.76%
Wikihow Koupaee and Wang [2018] English WikiHow 128543 0.25 1.88%
Yahoo Answers (Title-Answer) Zhang et al. [2015] English Yahoo 1198018 0.05 3.50%
Many downstream retrieval benchmarks Thakur et al. [2021], Xiao et al. [2023] are evaluated out-of-domain, where no training samples from the downstream sources are available. To avoid the overfitting of training sets, as is shown in the above Table 6, we slightly tune the epoch used in the dataset sampling, by following previous works in Ni et al. [2022], Wang et al. [2024b], Ma et al. [2025].
@@
Training set decontamination is essential for developing fair, unbiased, and robust retrieval systems. Existing training datasets are derived from multiple sources, whose training data have possible duplicates with the test sets. Following tDRO Ma et al. [2025], we perform strict training set decontamination with SimHash Manku et al. [2007] to ensure no training queries appear in the BeIR Thakur et al. [2021] and CMTEB Retrieval Xiao et al. [2023]. The original sizes, deduplicated sizes, and duplicate numbers as listed in the Table 7.
@@ A.6 Task instructions
As previously mentioned in Section 2.1, modern LLM-based retrievers Wang et al. [2024b], BehnamGhader et al. [2024] introduce task-specific instructions to format input queries, which enables higher retrieval capacities. These LLM-based instruction-tuned retrieval is similar to instruction-based supervised fine-tuning (SFT) in LLM generation Ouyang et al. [2022]. Following Mistral-E5 Wang et al. [2024b], we reuse most of its task instructions, and format the instruction and query in the following format:
@@
The detailed instructions are listed below, most of which are directly copied from the previous works E5-Mistral Wang et al. [2024b] and tDRO Ma et al. [2025] for reproducibility. Note that one training set may correspond to multiple instructions to avoid overfitting.
@@ A.7 Efficiency optimizations techniques
, where the vocabulary dimension $V$ is much larger than the hidden dimension $H$. For example, for Llama3.1-8b Touvron et al. [2023] model, $V=128256$ and $H=4096$, where $\lfloor V/H\rfloor=31$. This means that, given batch size B, sequence length T, hidden size H, vocabulary size V, the intermediate tensors with shape $[B,T,V]$ in the projection, ReLU, and Log are 31x larger than a dense tensor $[B,T,H]$, which consumes huge amounts of GPU memory during training and inference. However, only a tensor with shape $[B,V]$ is needed for the final aggregated representations. Thus, an optimization for computing the sparse vector is needed urgently.
@@
Proper optimizations are important. For example, we found it hard to reproduce quickly with some of the open-sourced baselines BehnamGhader et al. [2024], which is roughly 6x slower than our implementation when testing with the same size of LLM. Our experiments involve training with 23 train sets and inference with another 23 test sets, which requires proper optimizations urgently. We import other optimizations based on the implementations  [^4] Please refer to Ma et al. [2025]’s implementation on https://github.com/ma787639046/tdro. released by tDRO Ma et al. [2025]. These optimizations include 1) A distributed multi-node inference framework based on PyTorch RPC Li et al. [2023]; 2) Sequence packing Kundu et al. [2024] with Flash Attention 2 Dao [2023], which gives roughly 50% training speedup due to the removal of the pad tokens.; 3) Fused Kernels from Liger-Kernel Hsu et al. [2024], which gives roughly 10% speedup to both training and inferencing; 4) Multi-vector gradient accumulation support with GradCache Gao et al. [2021] for optionally enlarging batch sizes if necassary.

### v4

@@ 4 Ablations
Table 3: Symmetry Ablation settings (A1-A3). 1L-Trans denotes a single-layer Transformer.
@@
A3 1L-Trans Full 1L-Trans Full
Table 4: Ablations (A1-3) results (nDCG@10) of LightRetriever.
Ablations BeIR CMTEB-R
Llama-1b 48.7 58.7
A1: Both-side Light 34.9-13.8 40.1-18.6
A2: Encq use a MLP 37.5-11.2 41.3-17.4
Llama-8b 54.4 63.0
A3: Encq use 1L-Trans 50.1-4.3 54.4-8.6
@@ A.2.1 Text retrieval
Text retrieval builds on top of the vector space model [63], which encodes query and document representation vectors, and performs relevance search based on vector similarities.
@@
Dense retrieval [30] trains PLM-based encoders to condense queries and passages into low-dimensional dense vectors, then performs relevance search based on the Maximum Inner Product Search (MIPS) [52] algorithm. In recent years, dense retrievers, such as BGE [11] and E5 [71], have gained popularity for strong retrieval abilities across different tasks [51, 66], languages [84], and context granulaties [75, 53], owing to diverse training data [60], retrieval-customized pre-training [22, 76, 47], improved negative-mining [79], and so on. Recent SOTA retrievers [72, 5, 39] start to utilize LLMs [27, 68] as backbone encoders. With larger parameters and pre-training data, LLM-based retrievers enable significantly powerful retrieval capacities and domain adaptation abilities. However, LLMs also incur heavy online workloads when serving as query encoders. Improving the serving efficiency of LLM-based retrievers is still left for exploration.
@@
Traditional statistical sparse retrieval algorithms (e.g., TF-IDF [64], and BM25 [61]) do not involve learnable language models. Instead, they directly analyze lexical frequencies (e.g., TF) and importance (e.g., IDF) in a large vocabulary space. These lexicals are sparsely distributed and require building inverted indexes [46] for feasible similarity metric calculation, such as Apache Lucene [6] and Tantivy [59]. In recent years, sparse retrieval has started to work with pre-trained LMs (PLMs). BGE [11], SPLADE [21], SparTerm [3], and SPARTA [85] use customized XLM-RoBERTa [14] or BERT [19] as backbone encoders to learn term frequencies in an end-to-end manner and directly perform impact search based on the dot product of sparse vectors. Because of sparsity, these works also search with inverted indexes. Recent works [86] also try to incorporate LLMs with sparse retrieval. However, the online efficiency issue still exists.
@@
Hybrid retrieval is a technique that combines retrieval scores [73] or ranks [15] from multiple retrieval systems, such as dense [30] / sparse [61, 21] retrieval, and single [30] / multiple [32, 11] vector retrieval. It enables better retrieval abilities by interpolating results from multiple sources. Linear interpolation [73] of dense and sparse scores is used in our work.
@@ A.2.2 Improving inference efficiency
Previous works have tried to distill small query encoders from large encoders for better query inference efficiency. KALE [9] prunes the BERT-based query encoder and distills it during post-training. However, when the query encoder is pruned to one Transformers layer, its recall@20 on NQ [37] decreases by 25-30%. [74] proposes asymmetric dual-encoders by pruning BERT layers or using smaller BERT models. Its experiments show that a 2-layer BERT query encoder retains 92.5% of the full-sized performances via distillation and proper student initialization. USTAD [33] also proposes similar asymmetry and distillation strategies, which choose the 12-layer BERT as teachers and the 6-layer DistilBERT or 2-layer BERT-mini as students. Its experiments show that the distilled BERT-mini query encoder achieves 95-97% of the teacher’s performance because of better distillation.
However, these studies have primarily focused on small language models (SLMs), such as BERT [19], and overlooked feasible approaches for improving online inference efficiency of LLM-based retrievers, where such optimization is even more critical. Furthermore, these existing methods typically retain Transformer layers in the query encoders, which limits their efficiency gains. In addition, most of these works rely on SLMs and are trained on relatively narrow datasets, such as MS-MARCO [53], thus unable to compete with LLM-based retrievers.
@@
Previous works have also been dedicated to reducing index sizes for smaller disk space usage. Matryoshka representation learning (MRL) [36] supports adjusting the dense vector dimensions by directly taking the top-k dimension of the original vector as shrink vectors. It aligns and trains the corresponding multi-dimensional vectors in the end-to-end contrastive learning. For example, if $k\in\{64,128,256\}$, query-passage dense vectors with $\{64,128,256\}$-dims could be obtained together after one encoder forward operation. For LM-based sparse retrieval, the similarity scores are directly affected by the impacts of corresponding lexicals [21]. Thus, it naturally supports reducing the index sizes by retaining Top-k terms/lexicals. To explore the potential of controlling index sizes, our work also explores shrinking index sizes with dense MRL and sparse Top-k in the Section A.3, which also shows strong generalization capacities.
@@ A.3 Additional results about controls of vector dimension and sparsity
Retrieval systems usually involve indexing millions of documents into search databases, which consumes large amounts of disk space, especially for dense vectors. For example, indexing 8.8M MS-MARCO passages [53] with Llama-8b model consumes 134.92 GB for Faiss dense retrieval and 11.24 GB for Lucene sparse retrieval. Thus, it’s essential to explore the possibility of controlling index sizes by shrinking vectors. 1) For dense retrieval, we incorporate Matryoshka Representation Learning (MRL) [36] with LightRetriever to support flexible controls of dense vector dimensions. Given a dense vector $v^{den}$ with dimension $H$, MRL derives dimension-shrunk vectors by simply cutting out the Top-k dimensions $v^{den}[:k]$, then applies the same training loss as the original vectors. MRL trains a collection of Top-k dimensions together, enabling flexible choosing of a proper k. 2) For sparse retrieval, the sparse vector directly learns the term impact (i.e., lexical frequency) within a large sparse dimension. Thus, LightRetriever supports controlling sparsity by directly taking the Top-k terms, where the larger impact terms matter more.
@@ A.4 Additional ablation about the effect of auxiliary KL loss
Previous works for BERT-based asymmetric retrievers [74, 33] emphasize the importance of alignment between asymmetric students and full symmetric teachers. To explore the effect of such alignment, besides contrastive loss as the main loss function, we also explore to adopt the KL loss as an auxiliary to align the asymmetric similarity scores $S^{student}=v_{q}^{den}\cdot v_{d}^{den}$ to full-sized symmetric similarity scores $S^{teacher}=Enc_{q}(q)\cdot Enc_{d}(d)$,
@@ A.5 Training datasets
Following the settings of previous LLM-based retrievers [72, 5, 48], our work uses 23 training sets pre-built by tDRO [48] or Sentence Transformers Training Data [60], including 20 English and 3 Chinese datasets. Detailed dataset information is listed in Table 6 as follows.
@@
Amazon Review (2018) [54] English Amazon $999$ 0.01 0.58%
cMedQA2 [82] Chinese Chinese Medical 99936 1 5.84%
DuReader [58] Chinese Chinese Web Collections 86366 1 5.05%
Eli5 [20] English Reddit 325390 0.05 0.95%
Fever [67] English Wikipedia QA 109808 0.8 5.13%
FiQA [49] English Financial 5498 1 0.32%
GooQA Pairs [31] English Web Collections 3012347 0.025 4.40%
HotpotQA [81] English Wikipedia QA 85000 0.5 2.48%
MSMARCO [53] English Web Collections 502854 1 29.39%
NFCorpus [7] English Medical 2585 1 0.15%
NPR [45] English News 594376 0.05 1.74%
NQ [37] English Wikipedia QA 58800 1 3.44%
PQA Pairs [41] English Wikipedia QA 99999 0.5 2.92%
Quora Duplicates Triples [26] English Forum Duplicates 97011 0.5 2.83%
S2ORC Title-Abstract [43] English Semantic Scholar 99998 0.5 2.92%
SciFact [70] English S2ORC 806 1 0.05%
SPECTER [13] English Semantic Scholar 136642 0.25 2.00%
StackExchange Duplicates (Title-Body) [17] English Forum Duplicates 250516 0.25 3.66%
E5 Synthetic [4] English GPT Synthetic Data 224791 1 13.14%
T2Ranking [78] Chinese Chinese Web Collections 200362 0.5 5.86%
Trivia [29] English Wikipedia QA 60370 0.5 1.76%
Wikihow [34] English WikiHow 128543 0.25 1.88%
Yahoo Answers (Title-Answer) [83] English Yahoo 1198018 0.05 3.50%
Many downstream retrieval benchmarks [66, 77] are evaluated out-of-domain, where no training samples from the downstream sources are available. To avoid the overfitting of training sets, as is shown in the above Table 6, we slightly tune the epoch used in the dataset sampling, by following previous works in [55, 72, 48].
@@
Training set decontamination is essential for developing fair, unbiased, and robust retrieval systems. Existing training datasets are derived from multiple sources, whose training data have possible duplicates with the test sets. Following tDRO [48], we perform strict training set decontamination with SimHash [50] to ensure no training queries appear in the BeIR [66] and CMTEB Retrieval [77]. The original sizes, deduplicated sizes, and duplicate numbers as listed in the Table 7.
@@ A.6 Task instructions
As previously mentioned in Section 2.1, modern LLM-based retrievers [72, 5] introduce task-specific instructions to format input queries, which enables higher retrieval capacities. These LLM-based instruction-tuned retrieval is similar to instruction-based supervised fine-tuning (SFT) in LLM generation [56]. Following Mistral-E5 [72], we reuse most of its task instructions, and format the instruction and query in the following format:
@@
The detailed instructions are listed below, most of which are directly copied from the previous works E5-Mistral [72] and tDRO [48] for reproducibility. Note that one training set may correspond to multiple instructions to avoid overfitting.
@@ A.7 Efficiency optimizations techniques
, where the vocabulary dimension $V$ is much larger than the hidden dimension $H$. For example, for Llama3.1-8b [68] model, $V=128256$ and $H=4096$, where $\lfloor V/H\rfloor=31$. This means that, given batch size B, sequence length T, hidden size H, vocabulary size V, the intermediate tensors with shape $[B,T,V]$ in the projection, ReLU, and Log are 31x larger than a dense tensor $[B,T,H]$, which consumes huge amounts of GPU memory during training and inference. However, only a tensor with shape $[B,V]$ is needed for the final aggregated representations. Thus, an optimization for computing the sparse vector is needed urgently.
@@
Proper optimizations are important. For example, we found it hard to reproduce quickly with some of the open-sourced baselines [5], which is roughly 6x slower than our implementation when testing with the same size of LLM. Our experiments involve training with 23 train sets and inference with another 23 test sets, which requires proper optimizations urgently. We import other optimizations based on the implementations  [^3] Please refer to [48]’s implementation on https://github.com/ma787639046/tdro. released by tDRO [48]. These optimizations include 1) A distributed multi-node inference framework based on PyTorch RPC [42]; 2) Sequence packing [35] with Flash Attention 2 [18], which gives roughly 50% training speedup due to the removal of the pad tokens.; 3) Fused Kernels from Liger-Kernel [24], which gives roughly 10% speedup to both training and inferencing; 4) Multi-vector gradient accumulation support with GradCache [23] for optionally enlarging batch sizes if necassary.